### v1

# Inflation in light of ACT/SPT: a new perspective from Weyl gravity

Abstract
Recent measurements from the Atacama Cosmology Telescope (ACT) and the South Pole Telescope (SPT) have placed the strictest constraints on the primordial scalar perturbation spectrum, reporting a spectral index of $n_{s}\sim 0.967-0.98$ at 95% confidence level. This result indicates a stronger scale invariance of the scalar perturbation than earlier estimates, posing challenges for numerous inflation models. In this work, we propose an appealing inflationary scenario from the Weyl scale-invariant gravity theory dominated by the higher-order curvatures. Specifically, the exponential curvature extensions are introduced to suppress the mass divergence of the inflaton. We find such scenario naturally yields leading-order predictions of $n_{s}\simeq 1-3/(2N)\sim 0.97-0.975$ or $n_{s}\simeq 1-5/(3N)\sim 0.967-0.972$ for various models, in excellent agreement with the ACT/SPT constraints. This result builds a concrete bridge between theoretical and observational scale invariance, implying an enduring cosmic echo of the primordial symmetry.
Introduction — It is well established that the observations of cosmic microwave background (CMB) reveal an almost scale-invariant scalar perturbation spectrum, characterized by a spectral index $n_{s}$ slightly less than 1. Recently, the Atacama Cosmology Telescope (ACT) [1, 2] and the South Pole Telescope (SPT) [3] have released updated CMB measurements, indicating an $n_{s}$ value closer to 1 than previously reported in Planck 2018 [4]. Specifically, $n_{s}\sim 0.967-0.98$ in the 95% confidence interval. This suggests that the scale invariance of the scalar spectrum is even stronger than earlier estimates.
An important goal of modern cosmology is to understand the origin of this approximate scale invariance. Generally, inflation stands as the leading theoretical framework capable of generating such a nearly scale-invariant spectrum [5]. Over the years, numerous inflation models have been developed. However, with the increasing precision of CMB observations—particularly following the release of the ACT/SPT results—many of these models are now facing stringent tests. For instance, well-studied scenarios such as Starobinsky inflation [6], Higgs inflation [7], and $\alpha$-attractors [8] predict a value of $n_{s}$ that is now too low to be consistent with the latest CMB data, leaving them disfavored at the 95% confidence level. These developments urge a reassessment of the inflationary mechanism, especially its intrinsic connection to scale invariance.
An enlightening idea posits that inflation could emerge from a gravity theory with local scale symmetry, also known as Weyl scale invariance. This symmetry is first introduced by H. Weyl [9, 10], and now has been widely applied in particle physics[11, 12, 13, 14, 15, 16, 17], cosmology[18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44], and quantum gravity [45, 46]. In such a framework of inflation, the spontaneous breaking of the Weyl symmetry would naturally give rise to a nearly scale-invariant scalar spectrum, implying a profound connection between the theoretical and observational scale invariance.
With Weyl symmetry, if gravity is formulated in a purely geometric and ghost-free manner, the most general allowed action reduces to the quadratic curvature term $\hat{R}^{2}$, where $\hat{R}$ denotes the Ricci scalar in Weyl geometry. This model is equivalent to a cosmological constant in the Einstein frame [15], corresponding to a pure de Sitter spacetime and a completely scale-invariant scalar spectrum. While, once matter is introduced, such as a scalar field, more scale-invariant terms are allowed in the action. They modify the induced cosmological constant to an effective scalar potential $V(\Phi)$ with asymptotically flat region in certain cases. This scenario is capable of realizing a viable inflation model and bringing about a slight deviation from the exact scale invariance of the spectrum present in the purely geometric case.
In previous studies, a scale-invariant linear curvature term $\phi^{2}\hat{R}$ is often introduced alongside the quadratic term to realize the above scenario [18, 19, 20, 21, 22]. This model leads to a Higgs-like potential for inflation. However, its predicted value for $n_{s}$ is lower than that of the Starobinsky model, thus further exacerbating the tension with current observational constraint. In this work, we explore the role of higher-order curvature extensions—rather than a linear term—in departing the pure de Sitter state derived from $\hat{R}^{2}$ model and explaining the tiny deviation of scalar spectrum from scale invariance. Specifically, we consider exponential curvature extensions, which, by contrast with simple polynomial terms, significantly alleviate the mass divergence of inflaton. We calculate the cosmological observables for such models both analytically and numerically, confirming their consistency with the ACT/SPT results.
These conventions are adopted in this work: FLRW metric $\mathrm{d}s^{2}=-\mathrm{d}t^{2}+a^{2}(t)\mathrm{d}\mathbf{x}^{2}$, natural unit $\hbar=c=1$, and Planck mass $M_{P}\equiv 1/\sqrt{8\pi G}=2.435\times 10^{18}\mathrm{GeV}$.
Weyl gravity — We start with a brief introduction to the Weyl $F(\hat{R},\phi)$ gravity, which contains a scalar $\phi$ and a Weyl gauge field $W_{\mu}\equiv g_{W}w_{\mu}$ with $g_{W}$ the gauge coupling. The Lagrangian in the Jordon frame is given by
$\frac{\mathcal{L}_{J}}{\sqrt{-g}}=\frac{1}{2}F(\hat{R},\phi)-\frac{1}{4g_{W}^{2}}F_{\mu\nu}F^{\mu\nu}-\frac{\zeta}{2}D_{\mu}\phi D^{\mu}\phi,$ (1)
where $\zeta$ is a free parameter, $D_{\mu}=\partial_{\mu}-W_{\mu}$ is the covariant derivative for scalar, $F_{\mu\nu}\equiv\partial_{\mu}W_{\nu}-\partial_{\nu}W_{\mu}$ is the field strength tensor of $W_{\mu}$, $\hat{R}$ denotes the Weyl Ricci scalar derived from the Weyl scale-invariant connection
$\hat{\Gamma}^{\rho}_{\mu\nu}=\Gamma^{\rho}_{\mu\nu}+\left(W_{\mu}g^{\rho}_{\nu}+W_{\nu}g^{\rho}_{\mu}-W^{\rho}g_{\mu\nu}\right),$ (2)
and $F(\hat{R},\phi)$ is a combination of $\hat{R}$ and $\phi$ preserving the scale invariance of Lagrangian. We introduce the following Weyl transformations,
$g_{\mu\nu}\to g^{\prime}_{\mu\nu}$ $=f^{2}(x)g_{\mu\nu},$ (3)
$\phi\to\phi^{\prime}$ $=f^{-1}(x)\phi,$ (4)
$W_{\mu}\to W^{\prime}_{\mu}$ $=W_{\mu}-\partial_{\mu}\ln|f(x)|.$ (5)
One can derive $\hat{R}\to f^{-2}(x)\hat{R}$ and $\sqrt{-g}\to f^{4}(x)\sqrt{-g}$, then verify the invariance of Eqs. (1) and (2) under these transformations.
To obtain the Lagrangian in the Einstein frame, we introduce a dimension-2 auxiliary field $\chi$ obeying $\chi=\hat{R}$, and rewrite the curvature term as
$F(\hat{R},\phi)=F_{,\chi}(\chi,\phi)\left(\hat{R}-\chi\right)+F(\chi,\phi),$ (6)
where $F_{,\chi}(\chi,\phi)\equiv\frac{\partial F(\hat{R}\to\chi,\phi)}{\partial\chi}$. By setting up a specific Weyl transformation with $f(x)=\sqrt{F_{,\chi}(\chi,\phi)}/M_{P}$, which amounts to the gauge fixing condition
$F_{,\chi}(\chi,\phi)=M_{P}^{2},$ (7)
the Lagrangian transfers to the Einstein frame,
$\frac{\mathcal{L}_{E}}{\sqrt{-g}}=\frac{M_{P}^{2}}{2}\hat{R}-\frac{1}{4g_{W}^{2}}F_{\mu\nu}F^{\mu\nu}-\frac{\zeta}{2}D_{\mu}\phi D^{\mu}\phi-V(\phi),$ (8)
$V(\phi)=-\frac{1}{2}\left[F(\chi,\phi)-\chi M_{P}^{2}\right].$ (9)
Clearly, the scale symmetry breaks spontaneously and an effective scalar potential emerges, where $\chi$ is expressed as a function of $\phi$ by solving Eq. (7).
Now we translate the Lagrangian to a more familiar form. Compared to the conventional Ricci scalar $R$, one can easily verify that
$\hat{R}=R-6W_{\mu}W^{\mu}-\frac{6}{\sqrt{-g}}\partial_{\mu}\left(\sqrt{-g}W^{\mu}\right).$ (10)
The total derivative term here is generally removed from the Lagrangian due to its null surface integral. While the second term can be combined with the $\phi$ kinetic term and together rewritten as
$\frac{\zeta}{2}D_{\mu}$ $\phi D^{\mu}\phi+3M_{P}^{2}W_{\mu}W^{\mu}=$
$\frac{1}{2}\frac{6M_{P}^{2}\zeta}{6M_{P}^{2}+\zeta\phi^{2}}\partial_{\mu}\phi\partial^{\mu}\phi+\frac{1}{2}\left(6M_{P}^{2}+\zeta\phi^{2}\right)\overline{W}^{2},$ (11)
where we have redefined the Weyl gauge field as
$\overline{W}_{\mu}\equiv W_{\mu}-\frac{1}{2}\partial_{\mu}\ln\left|6M_{P}^{2}+\zeta\phi^{2}\right|\equiv g_{W}\overline{w}_{\mu}.$ (12)
Further redefining the scalar field
$\frac{\phi^{2}}{M_{P}^{2}}\equiv\begin{cases}\frac{6}{|\zeta|}\sinh^{2}\left(\frac{\pm\Phi}{\sqrt{6}M_{P}}\right),~\zeta>0,\\ \frac{6}{|\zeta|}\cosh^{2}\left(\frac{\pm\Phi}{\sqrt{6}M_{P}}\right),~\zeta<0,\end{cases}$ (13)
the mass of Weyl gauge field is expressed as
$m_{W}^{2}(\Phi)$ $=g_{W}^{2}\left(6M_{P}^{2}+\zeta\phi^{2}\right)$
$=\begin{cases}+6g_{W}^{2}M_{P}^{2}\cosh^{2}\left(\frac{\pm\Phi}{\sqrt{6}M_{P}}\right),~\zeta>0,\\ -6g_{W}^{2}M_{P}^{2}\sinh^{2}\left(\frac{\pm\Phi}{\sqrt{6}M_{P}}\right),~\zeta<0,\end{cases}$ (14)
and the Lagrangian in the Einstein frame finally reduces to a conventional form,
$\frac{\mathcal{L}_{E}}{\sqrt{-g}}=$ $\frac{M_{P}^{2}}{2}R-\frac{1}{2}\partial_{\mu}\Phi\partial^{\mu}\Phi-V(\Phi)$
$-\frac{1}{4g_{W}^{2}}\overline{F}_{\mu\nu}\overline{F}^{\mu\nu}-\frac{1}{2}m_{W}^{2}(\Phi)\overline{w}_{\mu}\overline{w}^{\mu}.$ (15)
As can be seen, it provides an inflaton candidate $\Phi$ and a vector dark matter candidate $\overline{w}_{\mu}$ coupled with $\Phi$. Since negative $\zeta$ leads to $m_{W}^{2}<0$, we only focus on the $\zeta>0$ case in this work.
Models — In this framework, the curvature term $F(\hat{R},\phi)$ determines the configuration of the effective potential $V(\Phi)$, which can be derived from Eqs. (7), (9), and (13), albeit the analytic expression may be unavailable for some complex cases. With these equations, one can readily verify the conclusions we mentioned earlier: the purely geometric case $F=\alpha\hat{R}^{2}$ is equivalent to a cosmological constant, and an additional scalar $\phi$ is able to make the model dynamical, leading to a slight deviation from the scale-invariant spectrum.
A straightforward attempt is to introduce a polynomial extension,
$F(\hat{R},\phi)=\alpha\hat{R}^{2}+\lambda_{n}\hat{R}^{n}\phi^{4-2n}.$ (16)
As noted earlier, the $n=1$ case, which includes the linear curvature term, is widely investigated in previous studies. This model reduces to the purely geometric case (flat potential) for $\phi\to 0$, where the scale-invariant spectrum forms. Inflation occurs during $\phi$ evolves from 0 to the potential minimum, which is $\phi=\pm 1$ in this model. The deviation of $\phi$ from 0 when horizon crossing determines the deviation of spectrum from scale invariance. Specifically, the effective potential is a Higgs-like potential,
$V(\Phi)=\frac{M_{P}^{4}}{8\alpha}\left[1-\frac{6}{|\zeta|}\sinh^{2}\left(\frac{\Phi}{\sqrt{6}M_{P}}\right)\right]^{2},~\zeta>0.$ (17)
It predicts a spectral index $n_{s}\simeq n_{s}^{*}-11N/\zeta^{2}$ [36], where $n_{s}^{*}$ denotes the prediction in Starobinksy model. This result, however, is now disfavored by ACT/SPT constraints at the 95% confidence level. Therefore, we have to consider other possibilities.
Actually, the linear curvature term in the Jordan frame is unnecessary, as any form of $F(\hat{R},\phi)$ can induce the linear curvature in the Einstein frame after gauge fixing. With this in mind, we ponder on certain higher-order extensions rather than the linear term in this work. One may try the $n\geq 3$ case in Eq. (16). It reduces to the purely geometric case when $\phi\to\infty$ and creates a potential minimum at $\phi=0$. In other words, it develops a ”pit” on the cosmological constant potential around $\phi=0$. Inflation occurs during the evolution of $\phi$ from a large value to 0, and meanwhile induces the deviation of spectrum from scale invariance.
This model, however, is plagued by serious problems in the vicinity of $\phi=0$. For the $n=3$ case, the asymptotic form of the effective potential, $V_{\Phi\to 0}\propto|\Phi|M_{P}^{3}-\Phi^{2}M_{P}^{2}$, indicates that the inflaton’s effective mass remains $m_{\Phi}^{2}=V^{\prime\prime}(\Phi)<0$ after inflation, resulting in a permanent tachyonic instability. For cases with $n\geq 4$, the potential takes the form $V_{\Phi\to 0}\propto|\Phi|^{2-\frac{2}{n-1}}$, which leads to a divergent effective mass, $m_{\Phi}^{2}\propto|\Phi|^{-\frac{2}{n-1}}$, as $|\Phi|$ decreases. Thus, when the inflaton field falls below a critical value $|\Phi_{c}|$, the effective mass surpasses the Planck scale, thereby rendering the effective description invalid.
To alleviate this problem, we propose that the $\hat{R}^{2}$ model with exponential extensions has better behavior in the vicinity of $\phi=0$ than the simple polynomial extensions. There are several example models:
$F(\hat{R},\phi)=\alpha\hat{R}^{2}e^{\beta\hat{R}/\phi^{2}},$ (18)
$F(\hat{R},\phi)=\alpha\phi^{2}\hat{R}\left(e^{\beta\hat{R}/\phi^{2}}-1\right),$ (19)
$F(\hat{R},\phi)=\alpha\hat{R}^{2}\cosh\left(\beta\hat{R}/\phi^{2}\right),$ (20)
$F(\hat{R},\phi)=\alpha\phi^{2}\hat{R}\sinh\left(\beta\hat{R}/\phi^{2}\right),$ (21)
where $\alpha$ and $\beta$ are positive parameters. Since $\beta$ couples exclusively to $\phi^{2}$ in the exponential term, it can be absorbed by $\zeta$ in Eq. (13) via parameter redefinition. Therefore, we will omit it in the following discussions. Through series expansion, one can confirm that these models reduce to the model (16) at the sub-leading-order. Specifically, models (18) and (19) correspond to the $n=3$ case. We name this type of model as the type-A model. Models (20) and (21) correspond to the $n=4$ case, which are classed as the type-B model.
Note that the simplification of exponential-extended models with Eq. (16) is only valid in the regime $\phi^{2}\gg R$. The behavior near $\phi=0$, particularly the divergence behavior, is presented in different manners. Since Eq. (7) in these models lacks explicit solutions, we construct an asymptotic solution in the limit $\phi\to 0$. To leading-order, all models yield the same estimate: $\chi_{\phi\to 0}\simeq 2\phi^{2}\ln\left|\frac{M_{P}}{\phi}\right|$. Substituting this into Eq. (9) gives
$V_{\phi\to 0}\simeq\phi^{2}M_{P}^{2}\ln\left|\frac{M_{P}}{\phi}\right|,$ (22)
$V^{\prime\prime}(\Phi)\simeq\frac{M_{P}^{2}}{\zeta}\left(2\ln\left|\frac{M_{P}}{\Phi}\right|+\ln\zeta-3\right).$ (23)
The second derivative is still divergent, but the degree of divergence has reduced to the logarithmic level. This means the invalid interval of $\Phi$ where $m_{\Phi}>M_{P}$ becomes extremely narrow for sufficiently large $\zeta$. Specifically, it can be estimated to $|\Phi|<e^{-\zeta/2}$ for $\zeta\gg 1$. In fact, $\zeta$ allows for a very large value up to $10^{10}$ (shown in the next section), thus the premise of our discussion is satisfied.
Since $\Phi$ represents only the classical background field and non-negligible quantum fluctuations $\delta\phi$ are always present, one can expect that the characteristic scale of $\delta\phi$ is far larger than the narrow classically-invalid interval derived above. In other words, the interval where the classical analysis breaks down can be shielded by quantum effects. Therefore, the exponential-extended $\hat{R}^{2}$ models significantly alleviate the divergence problem inherent in the polynomial-extended models.
Figure 1: Effective potentials of Weyl exponential-extended models with $\alpha=10^{11}$, $\zeta=10^{8}$ as an example.
By solving Eq. (7) numerically, the effective scalar potentials of above exponential-extended models are depicted in Fig. 1. As expected, the $\hat{R}^{2}$ term dominates for $\Phi\to\infty$, inducing a flat potential with scale-invariant scalar spectrum. As $\Phi$ evolves, the higher-order curvatures become significant, leading to a gradual reduction of $n_{s}$ from 1. The parameter $\zeta$ controls the width of pit in the potential: larger values of $\zeta$ result in a wider pit, while smaller values lead to a narrower one. Actually, almost any higher-order curvature extension is capable of inducing potentials like this, but to relieve the divergence problem, the exponential extensions are more reasonable.
Observables — Now we calculate the inflationary observables, mainly the spectral index $n_{s}$ and the tensor-to-scalar ratio $r$. They are expressed as
$n_{s}=1-6\epsilon_{V}+2\eta_{V},~r=16\epsilon_{V},$ (24)
where the slow-roll parameters
$\epsilon_{V}\equiv\frac{M_{P}^{2}}{2}\left[\frac{V^{\prime}(\Phi)}{V(\Phi)}\right]^{2},~\eta_{V}\equiv M_{P}^{2}\frac{V^{\prime\prime}(\Phi)}{V(\Phi)}.$ (25)
We first present an analytical approach. Although these exponential-extended models lack an analytical expression for $V(\phi)$, they can be approximated by the polynomial model within the slow-roll region. We have classified these models into type-A and type-B in the previous section, respectively corresponding to the $n=3$ and $n=4$ cases of model (16). Both two cases admit explicit expressions for $V(\phi)$, allowing for analytical treatment.
We elaborate in detail on the type-A model as an illustrative example, which is $F(\hat{R},\phi)=\alpha\hat{R}^{2}+\lambda_{3}\hat{R}^{3}/\phi^{2}$ to the sub-leading-order. Solving Eq. (7), we have
$\chi=\frac{\alpha\phi^{2}}{3\lambda_{3}}\left(-1+\sqrt{1+\frac{3\lambda_{3}M_{P}^{2}}{\alpha^{2}\phi^{2}}}\right).$ (26)
For simplicity, the approximation $\phi^{2}\sim\Phi^{2}/\zeta$ is adopted. This is applicable for inflaton variations $\Delta\Phi\lesssim\sqrt{6}M_{P}$ after horizon crossing, which, within the model under consideration, corresponds to $\zeta\lesssim 10^{9}$. With this approximation and a redefined parameter $\xi\equiv\alpha^{2}/\left(3\lambda_{3}\zeta\right)$, the effective potential is derived as
$V(\Phi)\simeq-\frac{\xi M_{P}^{2}\Phi^{2}}{2\alpha}+\frac{\xi^{2}\Phi^{4}}{3\alpha}\left[\left(1+\frac{M_{P}^{2}}{\xi\Phi^{2}}\right)^{3/2}-1\right].$ (27)
Further defining $u\equiv\sqrt{1+M_{P}^{2}/(\xi\Phi^{2})}$ for convenience, the slow-roll parameters are expressed as
$\epsilon_{V}=2\xi\left(u^{2}-1\right)\left(\frac{u-1}{2u+1}\right)^{2},$ (28)
$\eta_{V}=-6\xi\frac{\left(u^{2}-1\right)\left(u-1\right)}{u\left(2u+1\right)},$ (29)
and the $e$-folding number becomes
$N$ $\simeq\int^{\Phi^{*}}_{\sim 0}\frac{\mathrm{d}\Phi}{\sqrt{2\epsilon_{V}}M_{P}}$
$=\frac{1}{2\xi}\left[\frac{3}{16}\ln\left(\frac{u-1}{u+1}\right)+\frac{\left(3u-1\right)\left(u+2\right)}{8\left(u-1\right)^{2}\left(u+1\right)}\right]^{u^{*}}_{\infty},$ (30)
where $\Phi^{*}$ denotes the field value when horizon crossing. For $N\sim 50-60$ and $\zeta\lesssim 10^{9}$, actually, the first logarithmic term can be ignored due to its negligible contribution. Further defining $u^{*}\equiv 1+\delta$ with $\delta\ll 1$, this equation reduces to $N\simeq\left(6+11\delta\right)/\left(32\xi\delta^{2}\right)$, and we obtain the following approximate solution,
$u^{*}\simeq 1+\frac{11+\sqrt{121+768\xi N}}{64\xi N}.$ (31)
Substituting it into the slow-roll parameters and expanding to the sub-leading-order, the inflationary observables are finally derived as
$n_{s}\simeq 1-\frac{3}{2N}-\frac{3\sqrt{3}}{8\sqrt{\xi}N^{3/2}},$ (32)
$r\simeq\frac{\sqrt{3}}{3\sqrt{\xi}N^{3/2}}+\frac{23}{48\xi N^{2}}.$ (33)
Clearly, for the type-A model, the spectral index reads $n_{s}\simeq 1-3/(2N)\sim 0.97-0.975$ at the leading-order. The type-B model can also be treated analytically in principle. However, it requires massive approximations and cumbersome steps. Skipping tedious calculations, we give a leading-order result: $n_{s}\simeq 1-5/(3N)\sim 0.967-0.972$. As a reminder, the ACT/SPT $2\sigma$ bounds lies in $n_{s}\sim 0.967-0.98$. Thus, a minor higher-order correction to the $\hat{R}^{2}$ model suffices to tilt the scalar spectrum from exact scale invariance to a level that closely matches the latest constraints.
Figure 2: Predictions of spectral index $n_{s}$ combined with tensor-to-scalar ratio $r$ in the Weyl exponential-extended models with various $\zeta$. The blue, green, red, and yellow dotted lines are respectively for models (18), (19), (20), and (21). The violet area is the latest observational constraints given by the ACT collaboration [2].
We now present the numerical results, as shown in Fig. 2. Evidently, the theoretical predictions are in good agreement with the ACT constraints for small $\zeta$, specifically, for $\zeta\lesssim 10^{9}$ in type-A models, and for $\zeta\lesssim 10^{10}$ in type-B models. Actually, a small $\zeta$ should be favored because it enable the sub-Planck evolution of inflaton in the Einstein frame. As $\zeta$ decreases, the predicted $n_{s}$ converges to a fixed range, matching the leading-order analytical results derived previously. Moreover, the tensor-to-scalar ratio is strongly depressed, scaling as $r\propto\zeta$ and $r\propto\zeta^{2}$ for type-A and type-B models, respectively.
Discussion — A comparison with conventional $f(R)$ theories is instructive. In typical $f(R)$ inflation, the higher-order extensions are treated as corrections to the Starobinsky model [47, 48, 49, 50, 51, 52, 53, 54, 55]. From the perspective of phenomenology, the Starobinsky prediction, $n_{s}\simeq 1-2/N\sim 0.96-0.967$, lies slightly below the ACT/SPT preferred range. Positive higher-order curvature corrections would further diminish $n_{s}$, exacerbating the tension with observations. Conversely, negative corrections could raise $n_{s}$ into agreement. However, the parameter fine-tuning is required to some extent. Our models are in a different situation. The leading-order result of $n_{s}$ conforms directly to the ACT/SPT constraints, thereby avoiding fine-tuning. Moreover, from the perspective of theoretical completeness, inflation in the $f(R)$ framework with negative corrections no longer originates from a de Sitter state with constant curvature. The Ricci scalar $R$ is not convergent toward the beginning of inflation, and a critical value of $R_{c}$ (or corresponding $\Phi_{c}$ in the Einstein frame) emerges where $f(R)$ or $f^{\prime\prime}(R)$ turns negative. This necessitates a cut-off at $R_{c}$, rendering the theory UV incomplete [47, 54]. In contrast, our models preserve the UV completeness, which can be verified that $R_{\phi\to\infty}=M_{P}^{2}/(2\alpha)$, corresponding to a constant-curvature origin.
We proceed to a brief discussion of the dark matter candidate in our models, the Weyl gauge boson $\overline{w}_{\mu}$. The production mechanism of this particle in the Weyl $\phi^{2}\hat{R}+\hat{R}^{2}$ model with a Higgs-like potential (Eq. (17)) has been investigated in detail in our previous work [35]. The inflation models proposed in this work, however, leads to a different outcome. First, the mass $m_{W}$ of $\overline{w}_{\mu}$ depends on the inflaton $\Phi$ during inflation (see Eq. (Inflation in light of ACT/SPT: a new perspective from Weyl gravity)), affecting the evolution of its longitudinal mode. In the Higgs-like potential model, the increase of $|\Phi|$ and $m_{W}$ during inflation provides negative contributions to the frequency term $\omega_{k}^{2}$ of the longitudinal mode, thus promoting the gravitational production of $\overline{w}_{\mu}$. By contrast, for models in this work, $|\Phi|$ and $m_{W}$ decrease during inflation, leading to a suppression of the gravitational production. We estimate that the yield can be reduced by two orders of magnitude compared to the previous scenario.
Meanwhile, these models are able to realize small-field inflation, which significantly lowers the Hubble parameter $H_{e}$ at the end of inflation and further suppresses the dark matter yield. Given that the relic energy fraction of gravitationally produced vector dark matter (for $m\ll H_{\mathrm{reh}}$) follows $\Omega_{\mathrm{DM}}\propto H_{e}^{2}m^{1/2}$ [56, 57, 58], we estimate that the minimum mass of $\overline{w}_{\mu}$ required to match the observed $\Omega_{\mathrm{DM}}\sim 0.26$ increases by many orders of magnitude—from $10^{-4}$ eV reported in [35] up to the eV or even GeV scale. This opens up the possibility for $\overline{w}_{\mu}$ to produce observable signatures via the Higgs portal, to which it also couples through the covariant kinetic term in the Weyl framework.
Moreover, in the $\phi^{2}\hat{R}+\hat{R}^{2}$ model, the coupling of $\Phi$ and $\overline{w}_{\mu}$ generates $\Phi\overline{w}_{\mu}\overline{w}^{\mu}$ at the leading-order due to the non-zero vacuum expectation value of $\Phi$. This term allows inflaton to decay spontaneously into $\overline{w}_{\mu}$, which would result in excessive dark radiation if the reheating temperature is relatively low. In contrast, for models in this work, the leading-order coupling takes the form $\Phi^{2}\overline{w}_{\mu}\overline{w}^{\mu}$, which forbids such spontaneous decays, thereby avoiding the overproduction of dark radiation. A detailed analysis of these issues will be presented in future work.
Conclusion — In this work, we demonstrate a viable scenario that inflation arises from the higher-order curvature extensions, especially the divergence-suppressed extensions with exponential forms, on the basis of the Weyl scale-invariant $\hat{R}^{2}$ model. In these models, a ”pit” is developed on the cosmological constant potential in the Einstein frame, thereby deviating the scalar perturbation spectrum from exact scale invariance. Remarkably, such deviations align perfectly with the latest CMB observations across a broad parameter range. At the leading-order, the spectral index is given by $n_{s}\simeq 1-3/(2N)$ or $n_{s}\simeq 1-5/(3N)$ for various models, which lies comfortably within the 95% confidence interval reported by ACT/SPT. Consequently, a scale-invariant inflationary framework naturally explains the observed nearly scale-invariant spectrum, implying an enduring cosmic echo of the primordial symmetry.

### v2

@@
 This result indicates a stronger scale invariance of the scalar perturbation than earlier estimates, posing challenges for numerous inflation models.
-In this work, we propose an appealing inflationary scenario from the Weyl scale-invariant gravity theory dominated by the higher-order curvatures.
+In this work, we propose a novel inflationary scenario based on Weyl scale-invariant gravity, where the quadratic curvature establishes the scale invariance of the scalar spectrum, while the higher-order extensions imprint the observed slight deviation.
 Specifically, the exponential curvature extensions are introduced to suppress the mass divergence of the inflaton.
@@
 Introduction — It is well established that the observations of cosmic microwave background (CMB) reveal an almost scale-invariant scalar perturbation spectrum, characterized by a spectral index $n_{s}$ slightly less than 1.
-Recently, the Atacama Cosmology Telescope (ACT) [1, 2] and the South Pole Telescope (SPT) [3] have released updated CMB measurements, indicating an $n_{s}$ value closer to 1 than previously reported in Planck 2018 [4].
-Specifically, $n_{s}\sim 0.967-0.98$ in the 95% confidence interval.
+Recently, the Atacama Cosmology Telescope (ACT) [1, 2] and the South Pole Telescope (SPT) [3] have released updated CMB measurements, indicating an $n_{s}$ value closer to 1 than previously reported in Planck 2018 [4], especially after incorporating the latest baryon acoustic oscillation data from DESI [5].
+Specifically, they constrain $n_{s}\sim 0.967-0.98$ in the 95% confidence interval.
 This suggests that the scale invariance of the scalar spectrum is even stronger than earlier estimates.
 An important goal of modern cosmology is to understand the origin of this approximate scale invariance.
-Generally, inflation stands as the leading theoretical framework capable of generating such a nearly scale-invariant spectrum [5].
+Generally, inflation stands as a compelling theoretical framework capable of generating such a nearly scale-invariant spectrum [6].
 Over the years, numerous inflation models have been developed.
 However, with the increasing precision of CMB observations—particularly following the release of the ACT/SPT results—many of these models are now facing stringent tests.
-For instance, well-studied scenarios such as Starobinsky inflation [6], Higgs inflation [7], and $\alpha$-attractors [8] predict a value of $n_{s}$ that is now too low to be consistent with the latest CMB data, leaving them disfavored at the 95% confidence level.
+For instance, well-studied scenarios such as Starobinsky inflation [7], Higgs inflation [8], and $\alpha$-attractor T-models [9] predict a value of $n_{s}$ that is now too low to be consistent with the latest CMB data, leaving them disfavored at the 95% confidence level.
 These developments urge a reassessment of the inflationary mechanism, especially its intrinsic connection to scale invariance.
@@
 This symmetry is first introduced by H.
-Weyl [9, 10], and now has been widely applied in particle physics[11, 12, 13, 14, 15, 16, 17], cosmology[18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44], and quantum gravity [45, 46].
+Weyl [10, 11], and now has been widely applied in particle physics [12, 13, 14, 15, 16, 17, 18], cosmology [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46], and quantum gravity [47, 48].
 In such a framework of inflation, the spontaneous breaking of the Weyl symmetry would naturally give rise to a nearly scale-invariant scalar spectrum, implying a profound connection between the theoretical and observational scale invariance.
@@
 However, its predicted value for $n_{s}$ is lower than that of the Starobinsky model, thus further exacerbating the tension with current observational constraint.
-In this work, we explore the role of higher-order curvature extensions—rather than a linear term—in departing the pure de Sitter state derived from $\hat{R}^{2}$ model and explaining the tiny deviation of scalar spectrum from scale invariance.
+In this work, we explore the role of higher-order curvature extensions—rather than a linear term—in departing the pure de Sitter state derived from $\hat{R}^{2}$ model.
+Such extensions are not simple corrections to prior models, but rather lead to a brand new inflationary scenario.
 Specifically, we consider exponential curvature extensions, which, by contrast with simple polynomial terms, significantly alleviate the mass divergence of inflaton.
-We calculate the cosmological observables for such models both analytically and numerically, confirming their consistency with the ACT/SPT results.
+We calculate cosmological observables for such models and confirm their consistency with the ACT/SPT results.
 These conventions are adopted in this work: FLRW metric $\mathrm{d}s^{2}=-\mathrm{d}t^{2}+a^{2}(t)\mathrm{d}\mathbf{x}^{2}$, natural unit $\hbar=c=1$, and Planck mass $M_{P}\equiv 1/\sqrt{8\pi G}=2.435\times 10^{18}\mathrm{GeV}$.
@@
 and $F(\hat{R},\phi)$ is a combination of $\hat{R}$ and $\phi$ preserving the scale invariance of Lagrangian. We introduce the following Weyl transformations,
-$g_{\mu\nu}\to g^{\prime}_{\mu\nu}$ $=f^{2}(x)g_{\mu\nu},$ (3)
-$\phi\to\phi^{\prime}$ $=f^{-1}(x)\phi,$ (4)
+$g_{\mu\nu}\to g^{\prime}_{\mu\nu}=f^{2}(x)g_{\mu\nu},~\phi\to\phi^{\prime}=f^{-1}(x)\phi,$
 $W_{\mu}\to W^{\prime}_{\mu}=W_{\mu}-\partial_{\mu}\ln|f(x)|.$ (3)
@@
 Clearly, the scale symmetry breaks spontaneously and an effective scalar potential emerges, where $\chi$ is expressed as a function of $\phi$ by solving Eq. (5).
-Now we translate the Lagrangian to a more familiar form. Compared to the conventional Ricci scalar $R$, one can easily verify that
+Now we translate the Lagrangian to a more familiar form. Compared to the conventional Ricci scalar $R$, one can verify that
 $\hat{R}=R-6W_{\mu}W^{\mu}-\frac{6}{\sqrt{-g}}\partial_{\mu}\left(\sqrt{-g}W^{\mu}\right).$ (8)
@@
 Since negative $\zeta$ leads to $m_{W}^{2}<0$, we only focus on the $\zeta>0$ case in this work.
-Models — In this framework, the curvature term $F(\hat{R},\phi)$ determines the configuration of the effective potential $V(\Phi)$, which can be derived from Eqs.
+Models — In this framework, $F(\hat{R},\phi)$ determines the configuration of the effective potential $V(\Phi)$, which can be derived from Eqs.
 (5), (7), and (11), albeit the analytic expression may be unavailable for some complex cases.
-With these equations, one can readily verify the conclusions we mentioned earlier: the purely geometric case $F=\alpha\hat{R}^{2}$ is equivalent to a cosmological constant, and an additional scalar $\phi$ is able to make the model dynamical, leading to a slight deviation from the scale-invariant spectrum.
+With these equations, one can readily verify the conclusions we mentioned earlier: the purely geometric case $F=\alpha\hat{R}^{2}$ is equivalent to a cosmological constant, and an additional scalar $\phi$ is able to make the model dynamical, leading to an inflation model.
 A straightforward attempt is to introduce a polynomial extension,
@@
 Note that the simplification of exponential-extended models with Eq.
-(16) is only valid in the regime $\phi^{2}\gg R$.
+(14) is only valid in the regime $\phi^{2}\gg\hat{R}$.
 The behavior near $\phi=0$, particularly the divergence behavior, is presented in different manners.
@@
 This means the invalid interval of $\Phi$ where $m_{\Phi}>M_{P}$ becomes extremely narrow for sufficiently large $\zeta$.
-Specifically, it can be estimated to $|\Phi|<e^{-\zeta/2}$ for $\zeta\gg 1$.
+Specifically, it can be estimated to $|\Phi|<e^{-\zeta/2}M_{P}$ for $\zeta\gg 1$.
 In fact, $\zeta$ allows for a very large value up to $10^{10}$ (shown in the next section), thus the premise of our discussion is satisfied.
@@
 Further defining $u\equiv\sqrt{1+M_{P}^{2}/(\xi\Phi^{2})}$ for convenience, the slow-roll parameters are expressed as
-$\epsilon_{V}=2\xi\left(u^{2}-1\right)\left(\frac{u-1}{2u+1}\right)^{2},$ (28)
+$\epsilon_{V}=2\xi\frac{\left(u^{2}-1\right)\left(u-1\right)^{2}}{\left(2u+1\right)^{2}},$ (26)
 $\eta_{V}=-6\xi\frac{\left(u^{2}-1\right)\left(u-1\right)}{u\left(2u+1\right)},$ (27)
@@
 $n_{s}$ $\simeq 1-\frac{3}{2N}-\frac{3\sqrt{3}}{8\sqrt{\xi}N^{3/2}},$ (30)
-$r\simeq\frac{\sqrt{3}}{3\sqrt{\xi}N^{3/2}}+\frac{23}{48\xi N^{2}}.$ (33)
+$r$ $\simeq\frac{1}{\sqrt{3\xi}N^{3/2}}+\frac{23}{48\xi N^{2}}.$ (31)
 Clearly, for the type-A model, the spectral index reads $n_{s}\simeq 1-3/(2N)\sim 0.97-0.975$ at the leading-order.
@@
 Discussion — A comparison with conventional $f(R)$ theories is instructive.
-In typical $f(R)$ inflation, the higher-order extensions are treated as corrections to the Starobinsky model [47, 48, 49, 50, 51, 52, 53, 54, 55].
+In typical $f(R)$ inflation, the higher-order extensions are treated as corrections to the Starobinsky model [49, 50, 51, 52, 53, 54, 55, 56, 57, 58].
 From the perspective of phenomenology, the Starobinsky prediction, $n_{s}\simeq 1-2/N\sim 0.96-0.967$, lies slightly below the ACT/SPT preferred range.
@@
 Conclusion — In this work, we demonstrate a viable scenario that inflation arises from the higher-order curvature extensions, especially the divergence-suppressed extensions with exponential forms, on the basis of the Weyl scale-invariant $\hat{R}^{2}$ model.
-In these models, a ”pit” is developed on the cosmological constant potential in the Einstein frame, thereby deviating the scalar perturbation spectrum from exact scale invariance.
+In these models, the $\hat{R}^{2}$-induced cosmological constant serves as the origin of scale invariance for the scalar perturbation spectrum, while the higher-order curvatures bring about the observed slight deviation.
 Remarkably, such deviations align perfectly with the latest CMB observations across a broad parameter range.